OpenStudy (anonymous):

Worldwide Shipping Company charges (4x + 280) dollars for shipping a package that weighs x pounds from Miami to Denver. Friendly Shipping Company charges (3x + 125) dollars to ship the same package. Using your answer from #28, if a package weighs 259 pounds, how much can you save by shipping it via Friendly Shipping Company?

OpenStudy (anonymous):

414

OpenStudy (anonymous):

\[(4x+280)-(3x+125)=155+x \]\[155+259=\text{$\$$414} \]